### v1

@@ 2 Notation and Formal Definitions
We follow broadly the definitions and notations used in [Ami+22a]. Our setting is a network with a collection $N=\{0,\ldots,n-1\}$ of $n$ nodes. The input is an $n\times n$ demand matrix $D$ with nonnegative rational entries, where $D_{ij}$ specifies the amount of data that needs to be transferred from node $i$ to node $j$.
To define feasible solutions for the problems, we consider a particular network flow instance derived from the demand matrix $D$. The network/graph $G$ has vertex set $N\times\{0,\ldots,T\}$. The edge set of $G$ is partitioned in edges $E_{\mathrm{phys}}$ that we will label as physical, and edges $E_{\mathrm{virt}}$ that we will label as virtual. The virtual edges are of the form $(i,t)\to(i,t+1)$ for all $i\in N$ and $t\in\{0,\ldots,T-1\}$. The physical edges are of the form $(i,t)\to(j,t+1)$ for some $i\neq j\in N$. Furthermore, we require the set of physical edges $M_{t}$ at time $t$ to form a matching in the complete bipartite graph given by node set $(\{(i,t)\}_{i\in N},\{(i,t+1)\}_{i\in N})$, i.e., each node $(i,t)$ and each node $(i,t+1)$ is the end point of at most one edge in $M_{t}$, for all $i\in N$ and for all $t\in\{0,\ldots,T-1\}$. For all $i,j\in N$ and $t\in\{1,\ldots,T\}$, let $\mathcal{P}(i,j,t)$ be the set of all directed paths from vertex $(i,0)$ to vertex $(j,t)$ in $G$. We define $\mathcal{P}:=\cup_{i,j\in N,t\geq 1}\mathcal{P}(i,j,t)$. An indirect fractinoal routing is a flow $f:\mathcal{P}\to\mathbb{R}_{\geq 0}$. We say that the routing (or flow) $f$ is feasible for demands $D$ if it satisfies the following constraints:
Capacity Constraints
For each edge $e\in E_{\mathrm{phys}}$, $\sum_{P\in\mathcal{P}}f(P)\cdot\mathbf{1}[e\in P]\leq 1$.
Demand Satisfaction
@@
If a routing $f$ routes $\alpha$ units of flow along a path $P$ from $(i,0)$ to $(j,t)$ that contains a physical edge $(a,s)\rightarrow(b,s+1)$ then $\alpha$ units of data with source $i$ and destination $j$ are transmitted from node $a$ to node $b$ at time $s$.
A routing $f$ is a direct routing if it sends positive flow only over paths that contain at most one physical edge. A routing $f$ is integral if every node has at most one incoming physical edge with positive flow, and at most one outgoing physical edge with positive flow at any time step. Note that with our definitions, direct and integral routings are a subset of indirect and fractional routings. A direct routing can be viewed as a sequence of fractional or integer matchings $M_{0},\ldots,M_{T-1}$ in the complete bipartite graph between two copies of the node set $N$ that covers $D$.
We say that a routing $f$ for the network $G$ has a makespan of $T$. The completion time $t$ for a unit of data with source $i$ and destination $j$ is the smallest $t$ such that this data arrives at a node of the form $(j,t)$. Thus, the total completion time of a schedule is then
@@ 3 Fractional Matching, Indirect Routing, and Average Completion Time
In this section we prove Theorem˜1. For clarity of explanation, we prove the theorem for the total sum of completion times, which is equivalent to the average (up to a factor of $1/\sum_{i,j}D_{ij}$). We start with a description of the polynomial-time algorithm $\mathrm{ALG}$} that gives us Theorem˜1.
Algorithm $\bm{{\mathrm{ALG}}}$ Description:
At each time $t$, the fractional matching $M_{t}$ is an arbitrary maximal fractional matching of the residual demand matrix $D(t)$.
@@
Sender Bound Problem Statement:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$, it must be the case that $\sum_{i:s_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:r_{i}=v}p_{i}$ can be unbounded.
Receiver Bound Problem Statement:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$ it must be the case that $\sum_{i:r_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:s_{i}=v}p_{i}$ can be unbounded.
In the sender bound problem there is no bound on the rate that data can be transferred into a node, and in the receiver bound problem there is no bound on the rate that data can be transferred out of a node. Instead of limiting the amount of flow that can be transferred into or out of a node to 1, as it may seem natural, we limit it to $\frac{1}{4}$. This setting arises due to our dual fitting argument. Note that
@@
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})=\sum_{i,j\in N}D_{ij}\alpha^{S}_{ij}-\sum_{i,t}\beta^{S}_{it}+\sum_{i,j\in N}D_{ij}\alpha^{R}_{ij}-\sum_{j,t}\beta^{R}_{jt}\ .$
@@
${\mathrm{ALG}}$ $\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R}-D_{ij})\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R})=\sum_{i,j\in N}\left(D_{ij}\alpha_{ij}^{S}+D_{ij}\alpha_{ij}^{R}\right)\ .$
@@
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})$ $=\sum_{i,j}D_{ij}\alpha^{S}_{ij}-\sum_{i,t}\beta^{S}_{it}+\sum_{i,j}D_{ij}\alpha^{R}_{ij}-\sum_{j,t}\beta^{R}_{jt}$
@@
$4\cdot{\mathrm{OPT}}+4\cdot{\mathrm{OPT}}$ $\geq{\mathrm{OPT}}^{R}+{\mathrm{OPT}}^{S}\geq\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})\geq\frac{1}{2}\cdot{\mathrm{ALG}}\ .$
@@ 4 Integral Matching and Indirect Routing
We use a variant of Valiant’s hypercube design [VB81a]. Assume w.l.o.g. that $n$ is a power of $2$, and represent each node $u\in\{1,\ldots,n\}$ in binary. That is, $u=(u_{1},\ldots,u_{\log_{2}(n)})$ where $u_{i}\in\{0,1\}$ for all $i$. The connection schedule $\mathcal{M}$ will consist of exactly $\log_{2}(n)$ matchings. For each $i\in\{1,\ldots,\log_{2}(n)\}$, the $i$th matching connects nodes which match in all but the $i$th coordinate. That is,
@@
To prove the lower bound, note that in order to route uniform demands $B/n$ for any $B>0$, there must be some routing path between every source-destination pair $u,v$. This is only possible when the number of distinct paths which leave any fixed node $u$ is at least $n$. Given some makespan bound $L$, the number of distinct paths leaving a fixed node $u$ using exactly $1\leq d\leq L$ physical edges is equal to $L\choose d$}. Thus, the number of distinct paths leaving $u$ using at most $L$ physical edges is exactly $\sum_{d=1}^{L}{L\choose d}=2^{L}-1$. Therefore, it must be the case that $n\leq 2^{L}$, or equivalently, $L\geq\log_{2}(n)$. Using the same argument on source $u$ for reaching $n/2$ destinations, we can see the median completion time of demands send from $u$ is at least $\log_{2}(n)/2$, hence the average completion time of all demands is $\Omega(\log n)$. ∎
@@
Figure 2: The time-expanded routing scheme of a hypercube with dimension $d=2$ and $n=9$.
@@
Let $d=\left\lceil\log_{2}n/\log_{2}B\right\rceil$ and assume w.l.o.g. that $n^{1/d}$ is an integer. Represent nodes $u\in\{0,\ldots,n-1\}$ as $d$-tuples $u=(u_{1},\ldots,u_{d})$, where each $u_{i}\in\{0,\ldots,n^{1/d}-1\}$. For our connection schedule, we use a single period of the Elementary Basis connection schedule (Section 4.1 of [Ami+22a]), illustrated in Figure˜2 for dimension parameter $d=2$ and explained in detail below. We will use multiplicity of matchings $\left\lceil B/n^{1/d}\right\rceil$. For each index $i\in\{1,\ldots,d\}$ and each scale factor $s\in\{1,\ldots,n^{1/d}-1\}$, we define the matching $M_{t=(i,s)}$ which connects nodes $u$ to nodes $v$ which match $u$ in all but the $i$th coordinates, and which differ from $u$ in the $i$th coordinate by exactly $s(\mbox{mod }n^{1/d})$. That is,
@@
$\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{Bn}{Tn}=\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{B}{d\cdot(n^{1/d}-1)\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}=\frac{B}{n^{1/d}\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}\leq 1\ .$
@@
$T$ $=\left\lceil\frac{B}{n^{1/d}}\right\rceil\cdot dn^{1/d}\leq dn^{1/d}+dB=\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil\cdot\left(n^{1/\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil}+B\right)$
$\leq\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot\left(n^{\frac{\log_{2}(B)}{\log_{2}(n)}}+B\right)=\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot 2B\in\mathcal{O}\left(\frac{B\log_{2}(n)}{\log_{2}(B)}\right).$
@@
Fix any integral routing scheme $f$ and let $T$ be its makespan. For any node pair $u,v$, let $S_{u,v}$ denote the minimum number of physical edges used to route demands from $u$ to $v$, and let $h$ be the median of $S_{u,v}$ over all node pairs $u,v$. Let $d=\frac{1}{3}\log_{2}n/\log_{2}B$. We distinguish whether $h\leq d$ or $h\leq d$.
@@
Case 2: $h\leq d$. By the definition of $h$, for half of the node pairs $f$ routes some flow using at most $h$ physical edges. Thus, there must exists a sender $u^{*}$ that can route flow to at least $n/2$ receivers using at most $h$ physical edges each. Therefore, there must exist at least $n/2$ distinct paths from $u^{*}$ that reach those receivers within makespan $T$ using at most $h$ physical edges. We next compute how many such paths exist. There are $T\choose i$} distinct paths using exactly $i$ physical edges that obtain makespan $T$. Thus, there are $\sum_{i=1}^{h}{T\choose i}$ distinct paths with at most $h$ physical edges within makespan $T$. Thus, it must be that $n/2\leq\sum_{i=1}^{h}{T\choose i}$.
@@
$\sum_{i=1}^{h}{T\choose i}\leq 2{T\choose h}\leq 2\frac{T^{h}}{h!}\ ,$
@@
Using Stirling’s approximation $h!\geq\sqrt{2\pi h}\cdot(\frac{h}{e})^{h}$, we can derive $T\geq\frac{h}{e}n^{1/h}\cdot\left(\frac{\sqrt{2\pi h}}{4}\right)^{1/h}\geq\frac{h}{e}n^{1/h}\ ,$ because $\left(\frac{\sqrt{2\pi h}}{4}\right)^{1/h}$ converges to one from above as $h\to\infty$.
@@
This is a natural application of Valiant Load Balancing [VB81a], which is described at a high level below. To route $r$ demand between any node pair $u,v$, first traffic is routed from $u$ to all intermediate nodes $i$ in the network in equal shares $r/n$, using $f_{unif}$. Then, traffic is routed from intermediate nodes to its destination using $f_{unif}$. This allows us to treat routing any demand $D$ with row and column sums no more than $B$ as if it is two copies of uniform demand, requesting $B/n$ demand between node pairs. ∎

### v2

@@ 2 Notation and Formal Definitions
We follow broadly the definitions and notations used in [4]. Our setting is a network with a collection $N=\{0,\ldots,n-1\}$ of $n$ nodes. The input is an $n\times n$ demand matrix $D$ with nonnegative rational entries, where $D_{ij}$ specifies the amount of data that needs to be transferred from node $i$ to node $j$.
We define a particular flow network derived from the demand matrix $D$ that allows us to characterize feasible solutions for the problems we consider. The graph (or network) $G$ has vertex set $N\times\{0,\ldots,T\}$, for some $T\in\mathbb{N}$. The edge set of $G$ is partitioned into edges $E_{\mathrm{phys}}$ that we will label as physical, and edges $E_{\mathrm{virt}}$ that we will label as virtual. The physical edges are of the form $(i,t)\to(j,t+1)$ for all $i\neq j\in N$ and $t\in\{0,\ldots,T-1\}$. The virtual edges are of the form $(i,t)\to(i,t+1)$ for all $i\in N$ and $t\in\{0,\ldots,T-1\}$. For all $i,j\in N$ and $t\in\{1,\ldots,T\}$, let $\mathcal{P}(i,j,t)$ be the set of all directed paths from vertex $(i,0)$ to vertex $(j,t)$ in $G$. We define $\mathcal{P}:=\cup_{i,j\in N,t\geq 1}\mathcal{P}(i,j,t)$. An indirect fractional routing is a flow $f:\mathcal{P}\to\mathbb{R}_{\geq 0}$. We say that the routing (or flow) $f$ is feasible for demands $D$ if it satisfies the following constraints:
Capacity Constraints:
each vertex has a total outgoing flow over physical edges of at most $1$, that is, for each $t\in\{0,\ldots,T-1\}$ and $i\in N$,
$\sum_{j\in N\setminus\{i\}}\sum_{P\in\mathcal{P}}f(P)\cdot\mathbf{1}[((i,t),(j,t+1))\in P]\leq 1\ ,$
and each vertex has a total incoming flow over physical edges of at most $1$, that is, for each $t\in\{0,\ldots,T-1\}$ and $j\in N$,
$\sum_{i\in N\setminus\{j\}}\sum_{P\in\mathcal{P}}f(P)\cdot\mathbf{1}[((i,t),(j,t+1))\in P]\leq 1\ .$
Demand Satisfaction:
@@
If a routing $f$ routes $\alpha$ units of flow along a path $P$ from $(i,0)$ to $(j,t)$ that contains a physical edge $(a,s)\rightarrow(b,s+1)$, then $\alpha$ units of data with source $i$ and destination $j$ are transmitted from node $a$ to node $b$ at time $s$.
A routing $f$ is an integral routing if every vertex in $G$ has at most one incoming physical edge with positive flow, and at most one outgoing physical edge with positive flow; hence, for all $t\in\{1,\ldots,T\}$, the physical edges between vertex sets $\{(i,t)\}_{i\in N}$ and $\{(j,t+1)\}_{j\in N}$ with positive flow form a matching in $G$. A routing $f$ is a direct routing if it sends positive flow only over paths that contain at most one physical edge. A direct fractional routing and a direct integral routing can thus be viewed as a sequence of fractional or integral matchings $M_{0},\ldots,M_{T-1}$ in the complete bipartite graph between two copies of the node set $N$ that covers $D$, respectively. Note that with our definitions, direct routings are a subset of indirect routings, and integral routings are a subset of fractional routings.
We say that a feasible routing $f$ for the network $G$ has a makespan of $T$, if the last unit of data was delivered at time $T$. The completion time $t$ for a unit of data with source $i$ and destination $j$ is the smallest $t$ such that this data arrives at a node of the form $(j,t)$. Thus, the total completion time of a schedule is then
@@ 3 Fractional Matching, Indirect Routing, and Average Completion Time
In this section we prove Theorem˜1. For clarity of explanation, we prove the theorem for the total sum of completion times, which is equivalent to the average completion time (up to a factor of $1/\sum_{i,j}D_{ij}$). We start with a description of the polynomial-time algorithm $\mathrm{ALG}$} that gives us Theorem˜1.
Algorithm $\bm{{\mathrm{ALG}}}$:
At each time $t$, the fractional matching $M_{t}$ is an arbitrary maximal fractional matching of the residual demand matrix $D(t)$, where for all $i,j\in N$,
$D_{ij}(t):=D_{ij}-\sum_{t^{\prime}=0}^{t-1}(M_{t^{\prime}})_{ij}$
where $(M_{t^{\prime}})_{ij}$ denotes the amount of data on edge $(i,j)$ in $M_{t^{\prime}}$.
@@
Sender Bound Problem:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$, we require that $\sum_{i:s_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:r_{i}=v}p_{i}$ can be unbounded.
Receiver Bound Problem:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$, we require that $\sum_{i:r_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:s_{i}=v}p_{i}$ can be unbounded.
In the sender bound problem there is no bound on the rate that data can be transferred into a node, and in the receiver bound problem there is no bound on the rate that data can be transferred out of a node. Instead of limiting the amount of flow that can be transferred into or out of a node to one, as it may seem natural, we limit it to $\frac{1}{4}$. This setting arises due to technical reasons and makes our dual fitting arguments more convenient. Note that
@@
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})\\ =\sum_{i,j\in N}D_{ij}\alpha^{S}_{ij}-\sum_{i\in N}\sum_{t=0}^{T}\beta^{S}_{it}+\sum_{i,j\in N}D_{ij}\alpha^{R}_{ij}-\sum_{j\in N}\sum_{t=0}^{T}\beta^{R}_{jt}\ .$
@@
${\mathrm{ALG}}$ $\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R}-D_{ij})$
$\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R})$
$=\sum_{i,j\in N}\left(D_{ij}\alpha_{ij}^{S}+D_{ij}\alpha_{ij}^{R}\right)\ .$
@@
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})$
$=\sum_{i,j\in N}D_{ij}\alpha^{S}_{ij}-\sum_{i\in N}\sum_{t=0}^{T}\beta^{S}_{it}+\sum_{i,j\in N}D_{ij}\alpha^{R}_{ij}-\sum_{j\in N}\sum_{t=0}^{T}\beta^{R}_{jt}$
@@
$4\cdot{\mathrm{OPT}}+4\cdot{\mathrm{OPT}}$ $\geq{\mathrm{OPT}}^{R}+{\mathrm{OPT}}^{S}$
$\geq\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})\geq\frac{1}{2}\cdot{\mathrm{ALG}}\ .$
@@ 4 Integral Matching and Indirect Routing
We use a variant of Valiant’s hypercube design [41]. Assume w.l.o.g. that $n$ is a power of $2$, and represent each node $u\in\{1,\ldots,n\}$ in binary. That is, $u=(u_{1},\ldots,u_{\log_{2}(n)})$ where $u_{i}\in\{0,1\}$ for all $i$. The connection schedule $\mathcal{M}$ will consist of exactly $\log_{2}(n)$ matchings. For each $i\in\{1,\ldots,\log_{2}(n)\}$, the $i$th matching connects nodes which match in all but the $i$th coordinate. That is,
@@
To prove the lower bound, note that in order to route uniform demands $B/n$ for any $B>0$, there must be some routing path between every source-destination pair $u,v$. This is only possible when the number of distinct paths which leave any fixed node $u$ is at least $n$. Given some makespan bound $L$, the number of distinct paths leaving a fixed node $u$ using exactly $1\leq d\leq L$ physical edges is equal to $\binom{L}{d}$. Thus, the number of distinct paths leaving $u$ using at most $L$ physical edges is exactly $\sum_{d=1}^{L}\binom{L}{d}=2^{L}-1$. Therefore, it must be the case that $n\leq 2^{L}$, or equivalently, $L\geq\log_{2}(n)$. Using the same argument on source $u$ for reaching $n/2$ destinations, we can see the median completion time of demands sent from $u$ is at least $\log_{2}(n)/2$, hence the average completion time of all demands is $\Omega(\log n)$. ∎
@@
Figure 2: The time-expanded integral routing scheme of a hypercube with dimension $d=2$ and $n=9$.
@@
Let $d=\left\lceil\log_{2}n/\log_{2}B\right\rceil$ and assume w.l.o.g. that $n^{1/d}$ is an integer. Represent nodes $u\in\{0,\ldots,n-1\}$ as $d$-tuples $u=(u_{1},\ldots,u_{d})$, where each $u_{i}\in\{0,\ldots,n^{1/d}-1\}$. For our connection schedule, we use a single period of the Elementary Basis connection schedule (Section 4.1 of [4]), illustrated in Figure˜2 for dimension parameter $d=2$ and explained in detail below. We will use multiplicity of matchings $\left\lceil B/n^{1/d}\right\rceil$. For each index $i\in\{1,\ldots,d\}$ and each scale factor $s\in\{1,\ldots,n^{1/d}-1\}$, we define the matching $M_{t=(i,s)}$ which connects nodes $u$ to nodes $v$ which match $u$ in all but the $i$th coordinates, and which differ from $u$ in the $i$th coordinate by exactly $s(\mbox{mod }n^{1/d})$. That is,
@@
$\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{Bn}{Tn}$ $=\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{B}{d\cdot(n^{1/d}-1)\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}$
$=\frac{B}{n^{1/d}\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}\leq 1\ .$
@@
$T$ $=\left\lceil\frac{B}{n^{1/d}}\right\rceil\cdot dn^{1/d}$
$\leq dn^{1/d}+dB$
$=\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil\cdot\left(n^{1/\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil}+B\right)$
$\leq\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot\left(n^{\frac{\log_{2}(B)}{\log_{2}(n)}}+B\right)$
$=\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot 2B\in\mathcal{O}\left(\frac{B\log_{2}(n)}{\log_{2}(B)}\right).$
@@
Fix any integral routing scheme $f$ and let $T$ be its makespan. For any node pair $u,v$, let $S_{u,v}$ denote the minimum number of physical edges used to route demands from $u$ to $v$, and let $h$ be the median of $S_{u,v}$ over all node pairs $u,v$. Let $d=\frac{1}{3}\log_{2}n/\log_{2}B$. We distinguish whether $h\geq d$ or $h<d$.
@@
Case 2: $h<d$. By the definition of $h$, for half of the node pairs $f$ routes some flow using at most $h$ physical edges. Thus, there must exist a sender $u^{*}$ that can route flow to at least $n/2$ receivers using at most $h$ physical edges each. Therefore, there must exist at least $n/2$ distinct paths from $u^{*}$ that reach those receivers within makespan $T$ using at most $h$ physical edges. We next compute how many such paths exist. There are $\binom{T}{i}$ distinct paths using exactly $i$ physical edges that obtain makespan $T$. Thus, there are $\sum_{i=1}^{h}\binom{T}{i}$ distinct paths with at most $h$ physical edges within makespan $T$. Thus, it must be that $n/2\leq\sum_{i=1}^{h}\binom{T}{i}$.
@@
$\sum_{i=1}^{h}\binom{T}{i}\leq 2\binom{T}{h}\leq 2\frac{T^{h}}{h!}\ ,$
@@
Using Stirling’s approximation $h!\geq\sqrt{2\pi h}\cdot(\frac{h}{e})^{h}$, we can derive
$T\geq\frac{h}{e}n^{1/h}\cdot\biggl(\frac{\sqrt{2\pi h}}{4}\biggr)^{1/h}\geq\frac{h}{e}n^{1/h}\ ,$
because $(\sqrt{2\pi h}/{4})^{1/h}$ converges to $1$ from above as $h\to\infty$.
@@
This is a natural application of Valiant Load Balancing [41], which is described at a high level below. To route $r$ demand between any node pair $u,v$, first traffic is routed from $u$ to all intermediate nodes $i$ in the network in equal shares $r/n$, using $f_{unif}$. Then, traffic is routed from intermediate nodes to its destination using $f_{unif}$. This allows us to treat routing any demand $D$ with row and column sums no more than $B$ as if it is two copies of uniform demand, requesting $B/n$ demand between node pairs. ∎